OpenStudy (anonymous):

a culture initially contains 200 bacteria . if the number of bacteria doubles every 2 hours , how many bacteria will be in the culture at the end of 12 hours ?

OpenStudy (anonymous):

my calculation says 25,949 can someone pls confirms this.....

OpenStudy (anonymous):

It is exponential growth. First, how many division : 12/2 = 6 Second, how many bacteia from start : 200 Then 200*2⁶ = 12800

OpenStudy (anonymous):

correct william i made a mistake, u r right....

OpenStudy (anonymous):

General formula : \[x(t)=a\times b^{t/\tau}\] Giving bacteria quantity x at time t, with a divition in b bacteria each tau hour.

OpenStudy (anonymous):

@sierra14 pls use william's answer, his answer is right

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

pls give the medal to william as he got the right answer :))

OpenStudy (anonymous):

i gave it to both of you